cos u − cos 2u = 0
Answer by lwsshak3(11628)

For interval[0,2π]
cos u − cos 2u = 0
cos(u)-(cos^2(u)-sin^2(u))=0
cos(u)-cos^2(u)+sin^2(u)=0
cos(u)-cos^2(u)+1-cos^2(u)=0
2cos^2(u)-cos(u)-1=0
(2cos(u)+1)(cos(u)-1)=0
..
2cos(u)+1=0
cos(u)=-1/2
u=2π/3,4π/3
or
cos(u)+1=0
cos(u)=-1
u=π